OpenStudy (anonymous):

What are the exact solutions of x^2-5x-1=0 ?

mathslover (mathslover):

you know about quadratic equation

OpenStudy (anonymous):

Answer choices are X= 5 + squrt 29 / 2 X= -5 + squrt 29 / 2 X= 5 + squrt 21 / 2 X= -5 + squrt 21 / 2

OpenStudy (anonymous):

The plus signs are the weird one though that have the minus in the also u kno what I mean? Idk what they're called haha

mathslover (mathslover):

\[\large{ax^2+bx+c=0}\] roots of the equation : \[\large{x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}}\] here a 1 b = -5 and c =-1 can u do that now?

mathslover (mathslover):

solve this : \[\large{\frac{5\pm \sqrt{5^2-4(1)(-1)}}{2(1)}}\]

OpenStudy (anonymous):

5 + sqrt20/2

OpenStudy (anonymous):

I got that so far

mathslover (mathslover):